# AECBench: A Hierarchical Benchmark for Knowledge Evaluation of Large Language Models in the AEC Field

### v1

@@ 3.3 Data review
4 Evaluation of large language models (LLMs)
@@ 4.1 Models
Nine widely used LLMs were selected, encompassing both open-source and proprietary architectures. with QwQ-32B, DeepSeek-R1, and GPT o3-mini featuring specialized optimizations for enhanced reasoning capabilities. The selected models and their information are shown in Table 1.
@@
Moonshot-v1-128k [40] Moonshot AI Undisclosed w/o
GLM-4-Plus [41] Zhipu Undisclosed w/o
Qwen-Turbo [42] Alibaba Undisclosed w/o
QwQ-32B [43] Alibaba 32B w/
DeepSeek-V3 [44] DeepSeek 671B w/o
DeepSeek-R1 [45] DeepSeek 671B w/
GPT-4o [46] OpenAI Undisclosed w/o
GPT o3-mini [47] OpenAI Undisclosed w/
Hunyuan-TurboS [48] Tencent Undisclosed w/o
@@ 4.2 Experiment settings
The evaluation experiments were conducted on the OpenCompass platform [49], which enables automated answer extraction from model outputs and standardized computation of benchmarking metrics through its modular architecture. The experiments were conducted under a one-shot setting, which involves evaluating the model with a single example provided for each task, as shown in Figure 6. In this case, the model is expected to understand the task and generalize from this single example with minimal guidance. All experiments were primarily carried out on a server featuring an NVIDIA 4090 GPU (24GB) and Linux Ubuntu 22.04.3 LTS.
@@ 4.3.1 General observations
A systematic evaluation of nine LLMs was conducted through the evaluation platform. As detailed in Table 2, the task-specific performance metrics for each LLM are visualized in Figures 7 and 8. These figures also include horizontal lines representing the average performance for each level and sub-level. The test results clearly suggest a decline in LLMs’ capabilities across the five cognitive levels. A similar downward trend is also evident across the three sub-levels of the Knowledge Application level. Both major inter-level gaps and fine-grained intra-level differences were successfully identified, thereby validating the soundness and necessity of the hierarchical evaluation framework.
@@
  • The table shows the evaluation results of nine LLMs across twenty-three tasks. All multiple-choice questions are evaluated using accuracy as the metric. For the classification task (Task 5-1-1), the evaluation metric is F1. For Tasks 5-1-4 and 5-1-5, Soft-F1 is used as the evaluation metric, and F0.5 is used as the evaluation metric for Task 5-1-2. For evaluation tasks (Tasks 5-2-1 and 5-2-2), Kendall rank correlation coefficient ($\tau$) is used as the evaluation metric. Kendall’s $\tau$ is a statistic used to measure the ordinal association between two measured quantities. It ranges from -1 to +1, where +1 indicates perfect agreement between the two rankings, -1 indicates perfect disagreement, and a value near 0 indicates no correlation. In Tasks 5-3-1, 5-3-2, and 5-3-3, given that the DeepSeek-R1 model serves as the expert model for benchmark scoring, all corresponding performance metrics are annotated with an asterisk (*) to alert readers to potential evaluation bias risks. To provide a unified and comparable view of performance across all tasks, the results for all metrics have been normalized to a 0-100 scale. As all observed Kendall’s $\tau$ values were positive ([0, 1]), they were scaled by 100 (Score = $\tau$ × 100) for clarity and consistent presentation. The numbers in bold in the table represent the best-performing model for each task.
@@
The DeepSeek-R1 demonstrates a significant performance advantage over DeepSeek-V3, excelling in 19 of 23 designed tasks. This superiority is most pronounced in tasks at the Knowledge Reasoning and Knowledge Calculation level and the Evaluation and Creation sub-level. A direct comparison between the two models is justified by their virtually identical model architectures and parameter scales. This finding suggests that, for models of equivalent scale and architecture, reinforced training that specifically targets reasoning capabilities yields substantial enhancements in handling complex tasks within the AEC sector. For context, other control pairs from the same organizations, such as Qwen-32B vs. Qwen-Turbo and GPT-4o vs. GPT-3-mini, were considered. However, a lack of publicly available data on their internal architectures and parameters precludes a meaningful comparison, and their performance benchmarks do not present a discernible trend.
Among all the models, DeepSeek-R1 exhibits superior performance, standing out as the best-performing model in 14 out of 23 tasks and is validated as the automated judge. However, despite its strong and broad capabilities, its performance underscores a broader finding: even the most advanced LLMs still have substantial room for improvement in handling complex AEC-related tasks.
@@ 4.3.2 Specific observations
At the Knowledge Memorization level, all models performed exceptionally well, with scores universally exceeding 95% on terminology memorization (Task 1-2). This strongly suggests that a considerable amount of foundational AEC knowledge is already embedded within the parameters of current LLMs.
@@
At the Knowledge Understanding and Reasoning levels, a notable weakness was observed in tasks requiring the interpretation of tabular data (Tasks 2-3 and 3-3). The performance of the LLMs was notably deficient, even falling below that of the supposedly more advanced application tasks (for example, information extraction tasks). Conversely, on the building type inference task (Task 3-3), performance unexpectedly approached the performance of some Knowledge Memorization tasks (for example, code memorization & retrieval task). An analysis of Task 3-3 indicates that its correct option often exhibits high semantic similarity with the question stem, primarily because of shared keywords. Consequently, successfully completing this task may not strictly require inductive reasoning.
At the Knowledge Calculation level, the lower scores in architectural design calculation (Task 4-1) and engineering economics calculation (Task 4-4) are attributable to the multi-step analysis and complex computations these problems require, posing a greater challenge to the models.
@@
At the Knowledge Application level, performance on the sub-levels was highly variable. In the Analysis tasks, models worked well on document classification and proofreading tasks (Tasks 5-1-1 & 5-1-2) but struggled notably with compliance checking (Task 5-1-3) and brand compliance verification (Task 5-1-4). Lower scores in both Task 5-1-3 and Task 5-1-4 highlight a broader weakness in the ability to accurately compare key information, a core requirement for both tasks. This difficulty is then compounded by the unique demands of each: the former requires complex reasoning and calculation, while the latter necessitates the accurate parsing of tables in Markdown format. In contrast, performance on the information extraction task (Task 5-1-5) showcases the LLMs’ proficiency in parsing and extracting information from long-context domain documents.
@@
In the Evaluation tasks, the modest Kendall’s tau coefficients (failing to surpass the 0.7 threshold for strong correlation) are attributable to the profound difficulty of shifting from simple pattern matching to the kind of high-fidelity, evidence-based reasoning process. This process requires LLMs to interpret key evidence in criteria that may be paraphrased, expressed implicitly, or scattered across a long context, posing a significant challenge to the goal of substituting human experts in rubric-based evaluations. In the Creation tasks, all models failed to surpass the 60-point threshold. The uniformly limited performance is attributable to the complex demands of professional document generation. This process requires a multi-faceted synthesis of factual accuracy, logical coherence, and structural norms, posing a significant challenge to the models’ advanced generative capabilities.
@@ 5.1 Systematic bias of LLM-as-a-Judge approach
Two calibration methods (i.e., isotonic regression and piecewise linear regression) were applied in this paper to mitigate the observed scoring bias [50, 51]. As illustrated in Figure 10(b), both methods successfully shifted the model’s performance curve closer to the ideal y=x line. This improvement is quantitatively confirmed by a significant reduction in the mean absolute error (MAE) from an initial 2.947 to 1.926 (isotonic regression) and 2.015 (piecewise linear regression). These results demonstrate that both methods effectively reduce the systematic bias of LLM-as-a-Judge, leading to predictions that align more closely with human expert scores.
@@ 5.2 Performance decline in tasks related to code tables
Figure 11 presents model performance in two types of comparative tasks designed to specifically assess the model’s ability to understand and reason over knowledge in code tables. A significant and consistent performance degradation on these two tasks could be clearly observed. To investigate the reasons for this decline in consistency, this paper provides the model with the necessary knowledge sources in two distinct approaches. These two approaches enable LLMs to generate responses based on tabular content embedded in codes, as delineated in Figure 12. Approach #1 converts tabular content into natural language context through expert-crafted textual descriptions. Approach #2 utilizes models with multi-modal capabilities (GPT-4.1 in this paper) to automatically transform image-based tables into HTML formats, which is less intuitive for human interpretation but inherently machine-parsable [52]. Table 3 shows the performance of models on these two tasks. Both Approach #1 and Approach #2 achieved significant improvements over the baseline results (’Original’). The average accuracies on these two tasks improved from a baseline of 45.27% and 40.65% to 98.94% and 87.27% with Approach #1, and to 72.67% and 68.24% with Approach #2, respectively. These significant improvements also suggest that the decline stems from the model’s inability to properly encode knowledge from the code tables.
@@
Moreover, this paper compares two effective approaches for transforming knowledge within tables. After being processed by these approaches, the tables can not only serve as context for the model, as shown in Figure 12, but also have their knowledge converted into a pre-training dataset for the model. The superior performance of Approach #1 over Approach #2 can be attributed to Approach #2’s inability to correctly reproduce the complex layout of tables from building codes, particularly those with spanning cells. Even with advanced multi-modal architectures or dedicated parsing methods, lossless conversion of tabular content remains unattainable [53, 54, 55]. This inherent limitation underscores a fundamental trade-off in transforming tabular data for LLMs: while manual, expert-driven methods can achieve high fidelity, they lack scalability; conversely, automated approaches, though rapid and scalable, often struggle to preserve the complete accuracy and integrity of the original content.
@@ Acknowledgment
This project is supported by Shanghai Qi Zhi Institute (Grant No.SQZ202309) and the Scientific Research Project Fund of Arcplus Group PLC (24-1L-0142-Z). We extend our gratitude to every engineer and researcher who contributed to this dataset.
@@ A.5 Knowledge Application
Figure 30: The instruction and an example of Task 5-1-5 Information Extraction (continued on next page)
@@
(continued on next page)
@@
(continued on next page)

### v2

@@ 4.3 Data review
5 Evaluation of large language models
@@ 5.1 Models
Nine widely used LLMs were selected, encompassing both open-source and proprietary models, with QwQ-32B, DeepSeek-R1, and GPT o3-mini featuring specialized optimizations for enhanced reasoning capabilities. The selected models and their information are shown in Table 1.
@@
Moonshot-v1-128k [undefaj] Moonshot AI Undisclosed w/o
GLM-4-Plus [undefak] Zhipu Undisclosed w/o
Qwen-Turbo [undefal] Alibaba Undisclosed w/o
QwQ-32B [undefam] Alibaba 32B w/
DeepSeek-V3 [undefan] DeepSeek 671B w/o
DeepSeek-R1 [undefao] DeepSeek 671B w/
GPT-4o [undefap] OpenAI Undisclosed w/o
GPT o3-mini [undefaq] OpenAI Undisclosed w/
Hunyuan-TurboS [undefar] Tencent 560B w/o
@@ 5.2 Experiment settings
The evaluation experiments were conducted on the OpenCompass platform [undefas], which enables automated answer extraction from model outputs and standardized computation of benchmarking metrics through its modular architecture. The experiments were conducted under a one-shot setting, which involves evaluating the model with a single example provided for each task, as shown in Figure 7. In this case, the model is expected to understand the task and generalize from this single example with minimal guidance. While instruction-tuned models can perform well in zero-shot conditions [undefat, undefau], a single example is essential to unambiguously clarify task instructions and output formats across our 23 diverse tasks. More importantly, a few-shot setting was avoided because providing multiple demonstrations has been shown to "befuddle" advanced, instruction-tuned LLMs and, in some cases, even lead to a decline in model performance [undefav, undefaw, undefax]. Therefore, the one-shot setting provides the optimal balance: it offers the minimal guidance necessary for fair, unambiguous task execution, while ensuring testing the models’ internalized domain knowledge rather than their ability to over-fit to in-context examples. All experiments were primarily carried out on a server featuring an NVIDIA 4090 GPU (24GB) and Linux Ubuntu 22.04.3 LTS.
@@ 5.3.1 General observations
A systematic evaluation of nine LLMs was conducted through the evaluation platform. As detailed in Table 2, the task-specific performance metrics for each LLM are visualized in Figures 8 and 9. These figures also include horizontal lines representing the average performance for each level and sub-level. The test results clearly suggest a decline in LLMs’ capabilities across the five cognitive levels. A similar downward trend is also evident across the three sub-levels of the Knowledge Application level. Both major inter-level gaps and fine-grained intra-level differences were successfully identified, thereby validating the soundness and necessity of the hierarchical evaluation framework.
@@
  • The table shows the evaluation results of nine LLMs across twenty-three tasks. All multiple-choice questions are evaluated using accuracy as the metric. For the classification task (Task 5-1-1), the evaluation metric is F1. For Tasks 5-1-4 and 5-1-5, Soft-F1 is used as the evaluation metric, and F0.5 is used as the evaluation metric for Task 5-1-2. For evaluation tasks (Tasks 5-2-1 and 5-2-2), Kendall rank correlation coefficient ($\tau$) is used as the evaluation metric. Kendall’s $\tau$ is a statistic used to measure the ordinal association between two measured quantities. It ranges from -1 to +1, where +1 indicates perfect agreement between the two rankings, -1 indicates perfect disagreement, and a value near 0 indicates no correlation. In Tasks 5-3-1, 5-3-2, and 5-3-3, given that the DeepSeek-R1 model serves as the expert model for benchmark scoring, all corresponding performance metrics are annotated with an asterisk (*) to alert readers to potential evaluation bias risks. Though not explicitly shown here, self-enhancement bias [undefai] was analyzed and found to be marginal.To provide a unified and comparable view of performance across all tasks, the results for all metrics have been normalized to a 0-100 scale. As all observed Kendall’s $\tau$ values were positive ([0, 1]), they were scaled by 100 (Score = $\tau$ × 100) for clarity and consistent presentation. The numbers in bold in the table represent the best-performing model for each task.
@@
A direct comparison between the DeepSeek-R1 and DeepSeek-V3 is justified by their virtually identical model architectures and parameter scales. The DeepSeek-R1 demonstrates a significant performance advantage over DeepSeek-V3, excelling in 19 of 23 designed tasks. This superiority is most pronounced in tasks at the Knowledge Reasoning and Knowledge Calculation level and the Evaluation and Creation sub-level. This finding suggests that, for models of equivalent scale and architecture, reinforced training that specifically targets reasoning capabilities yields substantial enhancements in handling complex tasks within the AEC sector. For context, other control pairs from the same organizations, such as Qwen-32B vs. Qwen-Turbo and GPT-4o vs. GPT o3-mini, were considered. However, a lack of publicly available data on their internal architectures and parameters precludes a meaningful comparison, and their performance benchmarks do not present a discernible trend.
Among all the models, DeepSeek-R1 exhibits superior performance, standing out as the best-performing model in 14 out of 23 tasks and is validated as the automated judge. The top performance on the remaining nine tasks is dispersed among the other models, and DeepSeek-V3’s average performance is second only to DeepSeek-R1. Hunyuan-Turbo-S, the most recently released closed-source model, follows DeepSeek-V3 and outperforms DeepSeek-R1 on four specific tasks. Surprisingly, QwQ-32B, an open-source model with the smallest known parameter size among those tested, performs above average. Its performance is comparable to that of closed-source models such as Qwen-Turbo, GPT-4o, and GPT o3-mini. However, despite DeepSeek-R1’s strong and broad capabilities, its performance declines significantly, particularly on the tasks at the Knowledge Application level, a trend observed across all models. The result underscores a broader finding: even the most advanced LLMs still have substantial room for improvement in handling complex AEC-related tasks.
@@ 5.3.2 Specific observations
At the Knowledge Memorization level, all models performed exceptionally well, with scores universally exceeding 95% on terminology memorization (Task 1-2). Across these three tasks, model performance exhibited a high mean and low standard deviation. This strongly suggests that a considerable amount of foundational AEC knowledge is already embedded within the parameters of current LLMs.
@@
At the Knowledge Understanding and Reasoning levels, a notable weakness was observed in tasks requiring the interpretation of tabular data (Tasks 2-3 and 3-3). The performance of the LLMs was notably deficient, even falling below that of the supposedly more advanced application tasks (for example, information extraction tasks). The performance decline is discussed in detail in Section6. Conversely, on the building type inference task (Task 3-2), performance unexpectedly approached the performance of some Knowledge Memorization tasks (for example, code memorization & retrieval task). An analysis of Task 3-3 indicates that its correct option often exhibits high semantic similarity with the question stem, primarily because of shared keywords. Consequently, successfully completing this task may not strictly require inductive reasoning.
@@
At the Knowledge Calculation level, model performance across the four tasks exhibited a low mean and a high standard deviation. Particularly, the lower scores in architectural design calculation (Task 4-1) and engineering economics calculation (Task 4-4) are attributable to the multi-step analysis and complex computations these problems require, posing a greater challenge to the models. The high standard deviation precisely indicates that this capability is a key differentiator between high-performing models and general models. At the Knowledge Application level, performance on the sub-levels was highly variable. In the Analysis tasks, models worked well on document classification and proofreading tasks (Tasks 5-1-1 & 5-1-2) but struggled notably with compliance checking (Task 5-1-3) and brand compliance verification (Task 5-1-4). Lower scores in both Task 5-1-3 and Task 5-1-4 highlight a broader weakness in the ability to accurately compare key information, a core requirement for both tasks. This difficulty is then compounded by the unique demands of each: the former requires complex reasoning and calculation, while the latter necessitates the accurate parsing of tables in Markdown format. In contrast, performance on the information extraction task (Task 5-1-5) showcases the LLMs’ proficiency in parsing and extracting information from long-context domain documents, with the Moonshot-v1-128k model achieving the best performance on this task.
@@
In the Evaluation tasks, the modest Kendall’s tau coefficients (failing to surpass the 0.7 threshold for strong correlation) are attributable to the profound difficulty of shifting from simple pattern matching to the kind of high-fidelity, evidence-based reasoning process. This process requires LLMs to interpret key evidence in criteria that may be paraphrased, expressed implicitly, or scattered across a long context, posing a significant challenge to the goal of substituting human experts in rubric-based evaluations. The human inter-annotator agreement (IAA) baseline and the subsequent LLM-human consensus correlation are presented in Section 6. In the Creation tasks, all models failed to surpass the 60-point threshold. The uniformly limited performance is attributable to the complex demands of professional document generation. This process requires a multi-faceted synthesis of factual accuracy, logical coherence, and structural norms, posing a significant challenge to the models’ advanced generative capabilities.
@@ 6.1 Systematic bias of LLM-as-a-Judge approach
Two calibration methods (i.e., isotonic regression and piecewise linear regression) were applied in this paper to mitigate the observed scoring bias [undefay, undefaz]. As illustrated in Figure 11(b), both methods successfully shifted the model’s performance curve closer to the ideal y=x line. This improvement is quantitatively confirmed by a significant reduction in the mean absolute error (MAE) from an initial 2.947 to 1.926 (isotonic regression) and 2.015 (piecewise linear regression). These results demonstrate that both methods effectively reduce the systematic bias of LLM-as-a-Judge, leading to predictions that align more closely with human expert scores.
@@ 6.2 Performance decline in tasks related to code tables
Figure 12 presents model performance in two types of comparative tasks designed to specifically assess the model’s ability to understand and reason over knowledge in code tables. A significant and consistent performance degradation on these two tasks could be clearly observed. To investigate the reasons for this decline in consistency, this paper provides the model with the necessary knowledge sources in two distinct approaches. These two approaches enable LLMs to generate responses based on tabular content embedded in codes, as delineated in Figure 13. Approach #1 converts tabular content into natural language context through expert-crafted textual descriptions. Approach #2 utilizes models with multi-modal capabilities (GPT-4.1 in this paper) to automatically transform image-based tables into HTML formats, which is less intuitive for human interpretation but inherently machine-parsable [undefaaa]. Table 3 shows the performance of models on these two tasks. Both Approach #1 and Approach #2 achieved significant improvements over the baseline results (’Original’). The average accuracies on these two tasks improved from a baseline of 45.27% and 40.65% to 98.94% and 87.27% with Approach #1, and to 72.67% and 68.24% with Approach #2, respectively. These significant improvements also suggest that the decline stems from the model’s inability to properly encode knowledge from the code tables.
@@
Moreover, this paper compares two effective approaches for transforming knowledge within tables. After being processed by these approaches, the tables can not only serve as context for the model, as shown in Figure 13, but also have their knowledge converted into a pre-training dataset for the model. The superior performance of Approach #1 over Approach #2 can be attributed to Approach #2’s inability to correctly reproduce the complex layout of tables from building codes, particularly those with spanning cells. Even with advanced multi-modal architectures or dedicated parsing methods, lossless conversion of tabular content remains unattainable [undefaab, undefaac, undefaad]. This inherent limitation underscores a fundamental trade-off in transforming tabular data for LLMs: while manual, expert-driven methods can achieve high fidelity, they lack scalability; conversely, automated approaches, though rapid and scalable, often struggle to preserve the complete accuracy and integrity of the original content.
@@ Acknowledgment
This project is supported by Shanghai Qi Zhi Institute (Grant No.SQZ202309) and the Scientific Research Project Fund of Arcplus Group PLC (Grant No.24-1L-0142-Z). We extend our gratitude to every engineer and researcher who contributed to this dataset.
@@ A.5 Knowledge Application
Figure 31: The instruction and an example of Task 5-1-5 Information Extraction
(continued on thenext page)
@@
(continued on the next page)
@@
(continued on the next page)
@@
(continued on the next page)
Figure 35: The instruction and an example of Task 5-3-2 Specialized Report Generation(continued)